Monitors manufactured by TSI Electronics have life spans that have a normal distribution with a standard deviation of 1800 hours and a mean life span of 14,000 hours. If a monitor is selected at random, find the probability that the life span of the monitor will be more than12,920 hours. Round your answer to four decimal places.
Answer #1

  Solution :

Given that ,

mean = = 14000

standard deviation = = 1800

P(x >23920 ) = 1 - P(x < 23920)

= 1 - P[(x - ) / < (129200-14000) / 1800]

= 1 - P(z < -0.06)

Using z table,

= 1 -0.4761

=0.5239